OpenStudy (anonymous):

A bag has 2 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a green marble, do not replace it, and pick a yellow marble?

jimthompson5910 (jim_thompson5910):

how many green marbles are there

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

how many total

OpenStudy (anonymous):

8

jimthompson5910 (jim_thompson5910):

so if you were just picking one marble, then the probability of picking green is 2/8

jimthompson5910 (jim_thompson5910):

but we're picking 2 marbles and we are NOT putting the first one picked back in

jimthompson5910 (jim_thompson5910):

so after you pick the first one, the total drops from 8 to 7 So the probability of picking yellow on the second selection is 3/7 (since there are 3 yellow out of 7 total left)

jimthompson5910 (jim_thompson5910):

Multiply these two together to get (2/8)*(3/7) = ???

OpenStudy (anonymous):

so would it be 3/28

jimthompson5910 (jim_thompson5910):

(2/8)*(3/7) = (2*3)/(8*7) (2/8)*(3/7) = 6/56 (2/8)*(3/7) = 3/28 So you are 100% correct

jimthompson5910 (jim_thompson5910):

3/28 is the answer in fraction form